# Deepfakes, Phrenology, Surveillance, and More! A Taxonomy of AI Privacy Risks

### v1

@@ Abstract.
Privacy is a key principle for developing ethical AI technologies, but how does including AI technologies in products and services change privacy risks? We constructed a taxonomy of AI privacy risks by analyzing 321 documented AI privacy incidents. We codified how the unique capabilities and requirements of AI technologies described in those incidents generated new privacy risks, exacerbated known ones, or otherwise did not meaningfully alter the risk. We present 12 high-level privacy risks that AI technologies either newly created (e.g., exposure risks from deepfake pornography) or exacerbated (e.g., surveillance risks from collecting training data). One upshot of our work is that incorporating AI technologies into a product can alter the privacy risks it entails. Yet, current privacy-preserving AI/ML methods (e.g., federated learning, differential privacy) only address a subset of the privacy risks arising from the capabilities and data requirements of AI.
@@ 1. Introduction
These examples illuminate the unique privacy risks posed by AI technologies, prompting the foundational research question we ask in this work: How do modern advances in AI and ML change the privacy risks of a product or service? To answer this question, we introduce a taxonomy of AI privacy risks, grounded in an analysis of 321 privacy-relevant incidents that resulted from AI products and services, much like the ones described above. This work is important for at least two reasons. First, people are concerned about how AI can affect their privacy: a 2021 survey with around 10,000 participants from ten countries found that roughly half of the respondents believed that AI would result in “less privacy” in the future, citing concerns around large-scale collection of personal data, consent, and surveillance (Kelley et al., 2023). Second, while privacy is one of the five most commonly cited principles for the development of ethical AI technologies (Jobin et al., 2019), we do not yet have a systematic understanding of if and how modern advances in AI change the privacy risks entailed by products and services.
While AI and ML technologies have vastly expanded in capability (Yildirim et al., 2022), there is simultaneously a great deal of hype about what these technologies can and cannot do, making it difficult to separate real risks from speculative ones (Kapoor and Narayanan, Arvind, 2023). Thus, it can be difficult for today’s practitioners who develop AI-inclusive products and services to understand how their use of AI technologies might entail or exacerbate practical privacy risks (Yurrita et al., 2022). Prior work shows this to be true: in an interview with 35 AI practitioners, Lee et al. found that participants had relatively low awareness of privacy risks unique to or exacerbated by AI, and had little incentive to and support in addressing and mitigating these risks (Lee et al., 2024).
@@
We outline how this taxonomy can be used to create tools that help educate practitioners, and as a repository of shared knowledge regarding AI privacy risks and design processes to mitigate against those risks. Finally, we discuss how today’s conversations around privacy-preserving AI — which focus on approaches like differential privacy and federated machine learning — address only a subset of the risks we identify in our taxonomy. In sum, this paper makes the following contributions:
We codified a corpus of 321 documented AI incidents and present a taxonomy of privacy risks created and/or exacerbated by the capabilities and data requirement of AI technology.
Drawing from our taxonomy, we outline ways to aid practitioners in negotiating the utility-intrusiveness trade-off of AI-powered products and services with AI-specific privacy guidance.
@@ 2.1. Human-centered AI
To further develop what it means to design for privacy in human-centered AI, we present a novel taxonomy of AI privacy risks. This taxonomy aims to provide AI practitioners with tools and a shared language to foreground end-user privacy discussions in the design and development process.
@@ 2.2. Prior privacy taxonomies and concepts
It is crucial to turn to the literature on privacy law to address this gap and provide a more holistic and inclusive understanding of AI privacy risks as they manifest in the world. Legal discussions and the judicial system have addressed taxonomies of different types of privacy risks (Solove, 2006). Additionally, they have laid the groundwork of what truly qualifies as a “harm,” and how modern technology challenges these traditional distinctions (Citron et al., 2022). Nevertheless, these taxonomies do not possess the AI-specific focus of Shahriar et al.’s research. This paper intends to bridge the divide between these two bodies of work. By adopting an AI-centric approach to codify realized privacy risks, this paper introduces a distinctive taxonomy for evaluating and ultimately addressing privacy risks specific to the “life-cycle” of AI systems.
@@ 2.3. Creating a privacy taxonomy
The second characteristic that makes creating a universally accepted taxonomy difficult is the need to operationalize the taxonomy in a single domain. For example, there was a push for the taxonomy and approach of privacy by design for ubiquitous computing (Langheinrich, 2001). Similarly, in the field of Robotics, a specific taxonomy was developed to deal with implementing sensor technology (Eick and Anton, 2020). These taxonomies focus on the potential privacy risks the system or technology poses (Nunes and Brito, 2023). Even taxonomies built on user or societal input rely on perceived risks instead of reported harms resulting from past usage of similar technology (Jakobi et al., 2022). We identified one taxonomy that accounts for previously recorded privacy risks (Solove, 2006). However, this work is geared towards lawmakers and legal professionals rather than AI practitioners, and is not built to address specific privacy risks associated with the functionality and design of AI systems.
@@ 3. Method
We developed a taxonomy of privacy risks exhibited in documented AI privacy incidents by performing a systematic review of case studies. In this paper, we define AI broadly to accommodate the wide range of its capabilities to “perform tasks or behaviors that a person could reasonably deem to require intelligence if a human were to do it” (Riedl, 2019). AI is an umbrella term that encompasses many technologies, and our analysis does as well — we cover approaches ranging from Machine Learning (e.g., prediction and recommendation algorithms), Natural Language Processing (e.g., large language models), Computer Vision (e.g., facial recognition), and Robotics (e.g., home robots, drones). Note that we focus on documented end-user privacy risks of actual AI/ML products rather than speculative risks of general AI/ML concepts. We partly relied on the AI, Algorithmic, and Automation Incident and Controversy Repository (AIAAIC), the largest, up-to-date crowdsourced AI incident database curated by journalism professionals (Pownall, 2023). We also surveyed the AI Incident Database (AIID)111https://incidentdatabase.ai/, another public AI incident database, but decided not to use it because the AIAAIC provided good coverage of most of the privacy-related incidents in the AIID222We randomly selected 33% (N=50) of the AIID total incidents that contain the keyword “privacy” (N=151 as of August 16th, 2023). We manually went through the 50 incidents: 20 were either not AI products (e.g., augmented reality applications, executive orders from the government, policies) or not directly related to privacy (e.g., bias, accuracy), and 17 were already included in our AIAAIC database. The remaining 13 (26%) were not, but we found similar incidents in the AIAAIC database that were already captured by our taxonomy, e.g., incidents related to surveillance, data breaching, distortion made by deepfake AI, and physical invasion of AI technologies..
@@
Out of a database of 1,049 cases333We took a snapshot of the database on August 16th, 2023, 364 of them were labeled to involve “privacy issues”444The AIAAIC database tags each case with two attributes, Issue(s) and Transparency, to reflect if a given case raises any privacy concerns from the stakeholders and media. occurring between 2012 to 2023. We manually went through the linked resources for all 364 cases tagged in the AIAAIC as being privacy-pertinent and further filtered them down to 310 cases based on whether the reported technology (i) claimed to be inclusive of AI, ML, or otherwise “algorithmic” approaches, (ii) was actually deployed to real end-users, and (iii) involved some form of end-user privacy risks and/or compromise. We filtered out incidents that did not involve AI technologies (N=21, e.g., virtual-reality applications, data leakage unrelated to the use and development of AI), and incidents that were not associated with end-user privacy risks (N=33, e.g., bias (noa, 2021c), inaccuracy (Francisco, 2017), copyright (Gershgorn, 2021)).
@@
To ensure an adequate sampling strategy, we randomly picked 10% of the cases without the privacy label (69 out of 685) from the AIAAIC database. We identified privacy risk(s) in 11 of these cases (15.94%), and all of the identified risks were found in other cases tagged with the privacy label. Thus, we deemed our analysis had reached saturation. In sum, we analyzed a total of 321 distinct cases in developing our taxonomy of AI privacy risks (see Figure 2).
@@
We rooted our analysis on Solove’s taxonomy of privacy from 2006 as a baseline (Solove, 2006) — a popular conceptualization of privacy risks proposed prior to modern advances in AI/ML. Our primary analytic goal was to identify if and how AI exacerbates and/or creates privacy risks relative to this taxonomy, because doing so will highlight how modern advances in AI do and do not change notions of privacy risk. We say that AI exacerbates privacy risks when the capabilities and/or requirements of the AI technologies are not the root cause of the privacy risk, but increased its scale, scope, frequency, and/or intensity. We say that AI creates new privacy risks when the capabilities and/or requirements of the AI technology are fundamental enablers of the privacy risk. Otherwise, we say that the AI has not meaningfully changed the privacy risk described in the incident.
@@
For each incident, we assessed if and how the privacy violations described in the incident related to the unique context, capabilities of, and requirements entailed by the AI technologies described in the incident. We used an iterative coding process to categorize the privacy risk described in the incident. First, we created our codebook of different types of privacy risks adapted from the taxonomy proposed by Solove (Solove, 2006). Next, we iteratively updated the definition and scope of Solove’s initial set of privacy risks to be more specific to the AI privacy incidents in our dataset. For example, in our incident database, we observed that Increased Accessibility typically manifested as increasing public access to otherwise private or access-controlled data for building AI/ML models (e.g., through the release of public datasets). We also merged risks when they exclusively co-existed in our analysis. For example, we found that the Appropriation risk, the use of one’s identity to serve the aims and interests of another, always manifested with the Distortion risk, disseminating realistic AI-generated false information about individuals. While these two categories can theoretically be separable (i.e., one can imagine Distortion without Appropriation or Appropriation without Distortion), to keep our taxonomy grounded on real incidents and not theoretical harms, we merged the two categories into a single Distortion category. We also found an entirely new type of privacy risk, Phrenology / Physiognomy, which is not captured in Solove’s initial set of privacy risks. This privacy risk is unique to AI due to its capability to estimate sensitive personal attributes (e.g., sexual orientation, ethnicity) of individuals from their physical attributes (e.g., appearance, voice).
Finally, we excluded the “Breach of Confidentiality” risk in Solove’s taxonomy, as it was defined as an interpersonal risk between two people where one party discloses something to the other in confidence, and the other party violates this confidence by sharing it with third-parties. As AI technologies — even those that are designed to be anthropomorphic — are not capable of independently making promises, we consider this risk out of scope for AI technologies.
@@
In total, we created a final codebook of 12 operationalizable privacy risk labels for AI technologies, including Surveillance, Identification, Aggregation, Phrenology / Physiognomy, Secondary Use, Exclusion, Insecurity, Exposure, Distortion, Disclosure, Increased Accessibility, and Intrusion (see Table 2).
@@
To summarize our qualitative analysis procedure, the first author iteratively applied the codebook to 132 cases to update and better scope the definition of each privacy risk in active discussion with four other authors and constructed the initial codebook. Another author joined the coding process when the initial codebook was constructed. This author was trained with the codebook and independently coded the same set of 132 cases. The codes were then iteratively refined and discussed when disagreements occurred until both authors agreed on all codes in the codebook. To validate the inter-rater reliability, the two coders then independently coded another 65 cases (20% of our overall analysis pool; N=321) and reached a high agreement, with Cohen’s Kappa larger than 0.8 on every type of risk (see Table 1). One coder then coded the rest of the 124 cases. The final codebook comprises 12 types of privacy risks that we identified across the corpus of 321 cases. In determining whether AI newly created, exacerbated, or not meaningfully changed the privacy risks identified in each incident, the two coders engaged in an item-by-item discussion for all 321 incidents.
| Privacy Risk | Cohen’s Kappa |
| Surveillance | 0.88 |
| Identification | 0.94 |
| Secondary Use | 0.84 |
| Aggregation | 0.87 |
| Phrenology / Physiognomy | 1 |
| Exclusion | 0.90 |
| Insecurity | 0.94 |
| Exposure | 1 |
| Disclosure | 1 |
| Distortion | 1 |
| Increased Accessibility | 1 |
| Intrusion | 0.94 |
@@ 4. Taxonomy of AI Privacy Risks
Prime Voice AI, a text-to-voice generator, was misused to create the voices of celebrities to “make racist remarks about the US House representative”, and that the AI-generated clips “run the gamut from harmless, to violent, to transphobic, to homophobic, to racist” (Cox, 2023; Harrison, [n. d.]). |
@@ 4.1.1. Surveillance (150/321)
Scale (32/150): AI enhances the scale of surveillance in two ways: it enables linking across a diversity of sources and increases the quantity of collected personal data.
@@
Ubiquity (102/150): AI technologies can exacerbate surveillance risks by using physical sensors and devices to collect information from environments. For example, geolocation data from mobile devices were used to assess employee performance, raising concerns about employee tracking outside of work (Velazco, 2021). CCTV cameras have been used in applications to detect and prevent suicide attempts (Park, 2021) or to detect security anomalies in physical spaces (Wakefield, 2021b), while also introducing bystander privacy risks and concerns (Marky et al., 2020). Microphones enable a responsive audio interface for virtual assistants, along with concerns of extensive audio data collection and eavesdropping by the service provider (Paul, 2019).

### v2

@@ Abstract.
Privacy is a key principle for developing ethical AI technologies, but how does including AI technologies in products and services change privacy risks? We constructed a taxonomy of AI privacy risks by analyzing 321 documented AI privacy incidents. We codified how the unique capabilities and requirements of AI technologies described in those incidents generated new privacy risks, exacerbated known ones, or otherwise did not meaningfully alter the risk. We present 12 high-level privacy risks that AI technologies either newly created (e.g., exposure risks from deepfake pornography) or exacerbated (e.g., surveillance risks from collecting training data). One upshot of our work is that incorporating AI technologies into a product can alter the privacy risks it entails. Yet, current approaches to privacy-preserving AI/ML (e.g., federated learning, differential privacy, checklists) only address a subset of the privacy risks arising from the capabilities and data requirements of AI.
@@ 1. Introduction
These examples illuminate the unique privacy risks posed by AI technologies, prompting the foundational research question we ask in this work: How do modern advances in AI and ML change the privacy risks of a product or service? To answer this question, we introduce a taxonomy of AI privacy risks, grounded in an analysis of 321 privacy-relevant incidents that resulted from AI products and services, sourced from an AI incidents database (Pownall, 2023), much like the ones described above. This work is important for at least two reasons. First, people are concerned about how AI can affect their privacy: a 2021 survey with around 10,000 participants from ten countries found that roughly half of the respondents believed that AI would result in “less privacy” in the future, citing concerns around large-scale collection of personal data, consent, and surveillance (Kelley et al., 2023). Second, while privacy is one of the five most commonly cited principles for the development of ethical AI technologies (Jobin et al., 2019), we do not yet have a systematic understanding of if and how modern advances in AI change the privacy risks entailed by products and services.
While AI and ML technologies have vastly expanded in capability (Yildirim et al., 2022), there is simultaneously a great deal of hype about what these technologies can and cannot do, making it difficult to separate real risks from speculative ones (Kapoor and Narayanan, Arvind, 2023). Thus, it can be difficult for today’s practitioners who develop AI-inclusive products and services to understand how their use of AI technologies might entail or exacerbate practical privacy risks (Yurrita et al., 2022). Prior work shows this difficulty to be true: in an interview with 35 AI practitioners, Lee et al. found that participants had relatively low awareness of privacy risks unique to or exacerbated by AI, and had little incentive to and support in addressing these risks (Lee et al., 2024).
@@
We discuss how existing approaches to privacy-preserving AI and machine learning, such as differential privacy and federated machine learning, only account for a subset of these risks, highlighting the need for new tools, artifacts, and resources that aid practitioners in negotiating the utility-intrusiveness trade-off of AI-powered products and services. Finally, we outline how this taxonomy can be used to create tools that help educate practitioners, and as a repository of shared knowledge regarding AI privacy risks and design processes to mitigate against those risks.
@@ 2.1. Human-centered AI
We present a novel taxonomy of AI privacy risks to further develop what it means to design for privacy in human-centered AI. This taxonomy aims to provide AI practitioners with tools and a shared language to foreground end-user privacy discussions in the design and development process.
@@ 2.2. Prior privacy taxonomies and concepts
It is crucial to turn to the literature on privacy law to address this gap and provide a more holistic and inclusive understanding of AI privacy risks as they manifest worldwide. Solove’s work represents the progress within legal discussions and the judicial system to address taxonomies of different types of privacy risks (Solove, 2006). Solove’s taxonomy offered a comprehensive classification of different types of privacy intrusions (i.e., intrusions associated with information collection, information processing, information dissemination, and invasion) as seen in the legal field. It was previously used in security research to explore users’ personal attitudes and behaviors regarding privacy issues (Kokolakis, 2017; Acquisti et al., 2015). However, unlike this paper, the previous research did not look to apply or change the taxonomy to the new and emerging challenges of realized privacy intrusions.
Solove and colleagues built on their work by defining what qualifies as a “harm” and how modern technology challenges these traditional distinctions (Citron et al., 2022). Nevertheless, this taxonomy does not possess the AI-specific focus of Shahriar et al.’s research. Our paper intends to bridge the divide between these two bodies of work. By adopting an AI-centric approach to codify realized privacy risks, our paper introduces a distinctive taxonomy for evaluating and ultimately addressing privacy risks specific to the “life-cycle” of AI systems.
@@ 2.3. Creating a privacy taxonomy
The second characteristic that makes creating a universally accepted taxonomy difficult is the need to operationalize the taxonomy in a single domain. For example, there was a push for the taxonomy and approach of privacy by design for ubiquitous computing (Langheinrich, 2001). Similarly, in the field of Robotics, a specific taxonomy was developed to deal with implementing sensor technology (Eick and Anton, 2020). These taxonomies focus on the potential privacy risks the system or technology poses (Nunes and Brito, 2023). Even taxonomies built on user or societal input rely on perceived risks instead of reported harms resulting from past usage of similar technology (Jakobi et al., 2022). We identified one taxonomy that accounts for previously recorded privacy risks (Solove, 2006). However, this work is geared towards lawmakers and legal professionals rather than AI practitioners. It is not built to address specific privacy risks associated with the functionality and design of AI systems.
Similarly, previous HCI research has attempted to provide practitioners with a privacy taxonomy based on end users’ experience, raising awareness for physical privacy intrusions (Windl et al., 2023). This previous research shows the ability to apply such taxonomies in practical settings. Yet, it does not attempt to handle the more conceptual instances of AI privacy intrusions that may be invisible to end users but are no less impactful.
@@ 3. Method
3.1. Constructing the Taxonomy of AI Privacy Risks Based on AIAAIC
@@
We developed a taxonomy of privacy risks exhibited in documented AI privacy incidents by performing a systematic review of case studies. Creating typology and taxonomy by synthesizing real-world incidents has been used more broadly in privacy and security (Das et al., 2018) and in AI ethics (Raji et al., 2022). In this paper, we define AI broadly to accommodate the wide range of its capabilities to “perform tasks or behaviors that a person could reasonably deem to require intelligence if a human were to do it” (Riedl, 2019). AI is an umbrella term that encompasses many technologies, and our analysis does as well — we cover approaches ranging from Machine Learning (e.g., prediction and recommendation algorithms), Natural Language Processing (e.g., large language models), Computer Vision (e.g., facial recognition), and Robotics (e.g., home robots, drones). Note that we focus on documented end-user privacy risks of actual AI/ML products rather than speculative risks of general AI/ML concepts. We partly relied on the AI, Algorithmic, and Automation Incident and Controversy Repository (AIAAIC), the largest, up-to-date crowdsourced AI incident database curated by journalism professionals (Pownall, 2023). We also surveyed the AI Incident Database (AIID)111https://incidentdatabase.ai/, another public AI incident database, but decided not to use it because the AIAAIC provided good coverage of most of the privacy-related incidents in the AIID222We randomly selected 33% (N=50) of the AIID total incidents that contain the keyword “privacy” (N=151 as of August 16th, 2023). We manually went through the 50 incidents: 20 were either not AI products (e.g., augmented reality applications, executive orders from the government, policies) or not directly related to privacy (e.g., bias, accuracy), and 17 were already included in our AIAAIC database. The remaining 13 (26%) were not, but we found similar incidents in the AIAAIC database that were already captured by our taxonomy, e.g., incidents related to surveillance, data breaching, distortion made by deepfake AI, and physical invasion of AI technologies..
@@
Out of a database of 1,049 cases333We took a snapshot of the database on August 16th, 2023, 364 of them were labeled to involve “privacy issues”444The AIAAIC database tags each case with two attributes, Issue(s) and Transparency, to reflect if a given case raises any privacy concerns from the stakeholders and media. occurring between 2012 to 2023. To ensure that the incidents we analyzed indeed involve AI privacy risks, two coders reviewed the linked resources for all 364 cases tagged in the AIAAIC as being privacy-pertinent. Then, the two coders went through an incident-by-incident discussion on whether the reported technology (i) claimed to be inclusive of AI, ML, or otherwise “algorithmic” approaches, (ii) was actually deployed to real end-users, and (iii) involved some form of end-user privacy risks and/or compromise, and further filtered down to 310 cases. We filtered out incidents that did not involve AI technologies (N=21, e.g., virtual-reality applications, data leakage unrelated to the use and development of AI), and incidents that were not associated with end-user privacy risks (N=33, e.g., bias (noa, 2021c), inaccuracy (Francisco, 2017), copyright (Gershgorn, 2021)).
@@
To ensure an adequate sampling strategy, we randomly picked 10% of the cases without the privacy label in the AIAAIC database (69 out of 685). We identified privacy risk(s) in 11 of these cases (15.94%), and all of the identified risks were found in other cases tagged with the privacy label. Thus, we deemed our analysis had reached saturation. In sum, we analyzed a total of 321 distinct cases in developing our taxonomy of AI privacy risks (see Figure 2).
@@
As our objective was to understand how AI changes privacy, and not to re-define what is privacy, we rooted our analysis on Solove’s taxonomy of privacy from 2006 as a baseline (Solove, 2006) — a popular conceptualization of privacy risks proposed prior to modern advances in AI/ML. Our primary analytic goal was to identify if and how AI exacerbates and/or creates privacy risks relative to this taxonomy, because doing so will highlight how modern advances in AI do and do not change notions of privacy risk. We say that AI exacerbates privacy risks when the capabilities and/or requirements of the AI technologies are not the root cause of the privacy risk, but increased its scale, scope, frequency, and/or intensity — e.g., robust identification even with low-quality images. We say that AI creates new privacy risks when the capabilities and/or requirements of the AI technology are fundamental enablers of the privacy risk — e.g., deepfake pornography. Otherwise, we say that the AI has not meaningfully changed the privacy risk described in the incident.
@@
For each incident, we assessed if and how the privacy violations described in the incident related to the unique context, capabilities of, and requirements entailed by the AI technologies described in the incident. We used an iterative coding process to categorize the privacy risk described in the incident. First, we created our codebook of different types of privacy risks adapted from the taxonomy proposed by Solove (Solove, 2006). Next, we iteratively updated the definition and scope of Solove’s initial set of privacy risks to be more specific to the AI privacy incidents in our dataset. For example, in our incident database, we observed that Increased Accessibility typically manifested as increasing public access to otherwise private or access-controlled data for building AI/ML models (e.g., through the release of public datasets). We also merged risks when they exclusively co-existed in our analysis. For example, we found that the Appropriation risk, the use of one’s identity to serve the aims and interests of another, always manifested with the Distortion risk, disseminating realistic AI-generated false information about individuals. While these two categories can theoretically be separable (i.e., one can imagine Distortion without Appropriation or Appropriation without Distortion), to keep our taxonomy grounded on real incidents and not theoretical harms, we merged the two categories into a single Distortion category. Finally, we found an entirely new type of privacy risk, Phrenology / Physiognomy, which is not captured in Solove’s initial set of privacy risks. This privacy risk is unique to AI due to its capability to estimate sensitive personal attributes (e.g., sexual orientation, ethnicity) of individuals from their physical attributes (e.g., appearance, voice).
@@
In total, we created a final codebook of 12 operationalizable privacy risk labels for AI technologies, including Surveillance, Identification, Aggregation, Phrenology / Physiognomy, Secondary Use, Exclusion, Insecurity, Exposure, Distortion, Disclosure, Increased Accessibility, and Intrusion (see Table 1 and Figure 3).
3.2. Qualitative Analysis Procedure
@@
To summarize our qualitative analysis procedure, the first author iteratively applied the codebook to 132 cases to update and better scope the definition of each privacy risk in active discussion with four other authors and constructed the initial codebook. Another author joined the coding process when the initial codebook was constructed. This author was trained with the codebook and independently coded the same set of 132 cases. The codes were then iteratively refined and discussed when disagreements occurred until both authors agreed on all codes in the codebook. To validate the inter-rater reliability, the two coders then independently coded another 65 cases (20% of our overall analysis pool; N=321) and reached a high agreement, with Cohen’s Kappa larger than 0.8 on every type of risk and averaging 0.94 on all types of risks (see Appendix Table 2). One coder then coded the rest of the 124 cases. The final codebook comprises 12 types of privacy risks that we identified across the corpus of 321 cases. In determining whether AI newly created, exacerbated, or not meaningfully changed the privacy risks identified in each incident, the two coders engaged in an incident-by-incident discussion for all 321 incidents concerning the root cause of the privacy intrusions, as well as the role AI played in that root cause. The three themes — i.e., create, exacerbate, and no meaningful change — naturally emerged during this process.
@@ 4. Taxonomy of AI Privacy Risks
Prime Voice AI, a text-to-voice generator, was misused to create the voices of celebrities to “make racist remarks about the US House representative”, and that the AI-generated clips “run the gamut from harmless, to violent, to transphobic, to homophobic, to racist” (Cox, 2023; Harrison, 2023). |
@@ 4.1.1. Surveillance (150/321)
AI enhances the scale of surveillance (32/150) by enabling linking across a diversity of sources, and increasing the quantity of collected personal data.
@@
AI technologies exacerbate the ubiquity of surveillance risks (102/150) by using physical sensors and devices to collect information from environments. For example, geolocation data from mobile devices were used to assess employee performance, raising concerns about employee tracking outside of work (Velazco, 2021). CCTV cameras have been used in applications to detect and prevent suicide attempts (Park, 2021) or to detect security anomalies in physical spaces (Wakefield, 2021b), while also introducing bystander privacy risks and concerns (Marky et al., 2020). Microphones enable a responsive audio interface for virtual assistants, along with concerns of extensive audio data collection and eavesdropping by the service provider (Paul, 2019).